OpenStudy (anonymous):

Find the hypotenuse of a right triangle whose legs are sqrt of 10 and sqrt of 15. A. sqrt of 5 B. 5 C. 15

OpenStudy (welshfella):

hypotenuse^2 = h^2 = (sqrt10)^2 + (sqrt15)^2

OpenStudy (welshfella):

what is square root of 10 squared?

OpenStudy (welshfella):

if, for instance, i took the square root of 16 then squared the answer what would i get?

OpenStudy (welshfella):

or rhe square root of 4 and did the same thing

OpenStudy (welshfella):

what is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (welshfella):

right and 2 8 2 = ?

OpenStudy (welshfella):

2 * 2

OpenStudy (welshfella):

- come back to 4 right?

OpenStudy (anonymous):

ya

OpenStudy (welshfella):

so in the same way take sqrt 10 then i multiply sqrt10 by sqrt10 i get ?

OpenStudy (anonymous):

10 sqrt 10?

OpenStudy (welshfella):

sqrt4 * sqrt4 = 4 sqrt10 * sqrt10 = ?

OpenStudy (anonymous):

10

OpenStudy (welshfella):

yes1 by definition thats true also sqrt15 * sqrt15 = 15

OpenStudy (welshfella):

so:- hypotenuse^2 = h^2 = (sqrt10)^2 + (sqrt15)^2 h^2 = 10 + 15 h^2 = 25 so h = sqrt 25

OpenStudy (anonymous):

Oh! I got it now!!

OpenStudy (welshfella):

(remember (sqrt10)^2 = sqrt10 * sqrt10)

OpenStudy (welshfella):

good